OpenStudy (thomas5267):

If \(3x^2+4\) is an antiderivative of \(f(x)\), evaluate the indefinite integral \[F(x)=\int f(x)\,dx\] A: \(F(x)=x^3+4x+C\) B: \(F(x) = 3x^2+C\) C: \(F(x) = 3(x + C)^2+4\) D: \(F(x) = C(3x^2+4)\) E: none of the above Isn't the answer \(3x^2+4\)? That just seems wrong!

OpenStudy (callisto):

\[\int x^{n}dx = \frac{1}{n+1}x^{n+1}+C\]

OpenStudy (thomas5267):

OK, but the question is asking for the antiderivative of \(f(x)\), isn't that equals to \(\int f(x) \, dx\)? And the question says that it is \(3x^2+4\). WHAT!? The question has given you the answer!?

OpenStudy (callisto):

lol I see your point!!

OpenStudy (callisto):

Is it a quiz? homework question?

OpenStudy (thomas5267):

This is a assignment. If I am correct, \(F(x)=3x^2+4\). So does it equals to \(3x^2+C\)? But it is absolutely weird to use \(C\) in this context. The homework says that "Assume that C is an arbitrary constant throughout this assignment."

OpenStudy (callisto):

If \(3x^2+4\) is an antiderivative of f(x), so is \(3x^2+5\), \(3x^2+100\), etc. So, generally, \(3x^2+C\) is the antiderivative of f(x).

OpenStudy (callisto):

Yay! @mukushla comes to save us :P

OpenStudy (thomas5267):

OK. I get the point. \(3x^2+4\) is AN antiderivative of \(f(x)\), i.e. \(3x^2+4\) is one of the antiderivative of \(f(x)\). Therefore, the answer is B! *facepalm* Is this a question on Maths or English?

OpenStudy (callisto):

English :S and Maths :S

OpenStudy (callisto):

I'm sorry, I'm too careless!!

OpenStudy (thomas5267):

WTF? Playing with words in Maths assignment? YOU WIN! *ragequit*

OpenStudy (callisto):

Language :S